OpenStudy (anonymous):

How would I solve Integral of 7^(2x) dx? The indefinite Integral

OpenStudy (anonymous):

your question is?\[\int\limits_{}7^{2x}dx{}\]

OpenStudy (zzr0ck3r):

\(7^{2x}=e^{\ln(7^{2x})} =e^{2x*\ln(7)}\\\int7^{2x}dx=\int e^{2x*\ln(7)}dx=\\\frac{e^{2x}}{2\ln(7)}+c=\frac{49^x}{\ln(49)}+c\)

OpenStudy (zzr0ck3r):

that should say \(\frac{7^{2x}}{2\ln(7)}\) not \(\frac{e^{2x}}{2\ln(7)}\)

OpenStudy (anonymous):

thanks mate